# Rethinking On-policy Optimization for Query Augmentation

### v1

@@ 4.5 Discussions
5 On-policy Pseudo-document Query Expansion
5.1 Method
We make a simple modification to the on-policy optimization RL process. Instead of instructing the policy model to rewrite the query $q\rightarrow q^{\prime}$, we instruct the model to write a hypothetical document $d^{H}$, then directly output the concatenated $(q,d^{H})$ (prompt template in Appendix Appx.˜E). Then this concatenated string is used to retrieve relevant documents from the environment, similar to SPQE, and the reward is computed by the retrieval metrics like NDCG. In this way, we merge the SPQE into the on-policy optimization process. We refer to this method as On-policy Pseudo-document Query Expansion (OPQE).
5.2 Experiment, Result and Analysis
Experimental Setup. We experiment on evidence-seeking retrieval and ad hoc retrieval datasets. Our implementation is the same as Section˜4.2 except for the prompt template to instruct the policy model to write a pseudo document.
Main Results. We report the results of our hybrid approach in Table˜3. The key finding is that our proposed method, OPQE, successfully integrates the strengths of both prompting and on-policy optimization, achieving the best overall performance, especially in the dense retrieval setting.
@@ 5.2 Experiment, Result and Analysis
For dense retriever, the benefits of our hybrid approach are most apparent. On the ad hoc retrieval results, the OPQE-7B model achieves a new state-of-the-art average score of 58.1. This surpasses the previous best from the standard +7B RL model (57.5) and the prompt-based SPQE (56.6). This improvement is consistent across multiple datasets, with OPQE setting new top scores on benchmarks like DL20 (70.5). Similarly, on the evidence-based retrieval datasets, the OPQE-3B model pushes the average performance to 80.2, edging out all other methods on what is already a very strong baseline. This shows that by using RL to fine-tune the generation of pseudo-documents, our model learns to create query representations that are more effective for dense retrievers than either method could achieve alone.
@@
For the sparse BM25 retriever, the results demonstrate that OPQE is highly competitive. While the zero-shot SPQE remains the top performer on evidence-based datasets (78.2), our OPQE-7B model closes the gap significantly, outperforming the standard +7B RL model by a clear margin (76.5 vs. 75.6). On the Ad Hoc Retrieval Datasets, the OPQE-7B model matches the top performance of SPQE, both achieving an average score of 51.6. In summary, by framing on-policy optimization as generating a pseudo-document rather than a rewritten query, our OPQE method combines the robust generative structure of prompting with the targeted fine-tuning of reinforcement learning. This synergy leads to significantly improved performance, validating our approach as a superior method for query augmentation.
@@
Training Trajectories of OPQE. To better understand the performance gains of our hybrid OPQE method, we analyze its reward trajectories during training compared to the standard RL approach in Fig.˜1. These plots provide evidence for our hypotheses about the effectiveness of incorporating a pseudo-document generation structure.
@@
First, the trajectories validate our hypothesis that this structure effectively leverages the LLM’s internal knowledge, particularly for sparse retrievers. In Fig.˜1(a) and Fig.˜1(c) (FEVER and TriviaQA with BM25), the OPQE policy (blue line) starts with a significantly higher average reward than the standard RL policy (red line). This demonstrates that the initial, zero-shot pseudo-documents are already high-quality query expansions that provide rich lexical signals for the term-based BM25 retriever. OPQE essentially begins with a "warm start" by using the LLM’s knowledge, while the standard RL method must learn to generate effective query terms from a lower baseline.
@@ 6 Conclusion and Future Works
In this paper, we systematically evaluated and compared two LLM-based query augmentation paradigms: simple prompting-based query augmentation and RL-based query augmentation, across a variety of IR tasks. Our results suggest that prompt-based methods with advanced LLMs can perform on par with, or even surpass, RL-based methods in many scenarios, while RL-based approaches are particularly effective at enabling smaller LLMs to excel at this task. Furthermore, a hybrid approach, where pseudo-document generation prompts from prompt-based methods are used to provide stronger initialization for RL-based training, achieved the best overall performance. These findings suggest that RL-based query augmentation frameworks can be further optimized through improved task reformulation.
While our study highlights the strengths and trade-offs of prompting-based and RL-based query augmentation, several avenues remain open for future exploration. First, developing more lightweight yet effective prompting strategies for zero-shot query augmentation could reduce reliance on large LLMs and broaden applicability in resource-constrained settings. Second, refining RL task formulations, for example, by designing more nuanced reward functions, may help stabilize training and yield stronger performance. Finally, exploring query augmentation prompt design and RL training beyond the classic single-round retrieval paradigm, such as in multi-hop or deep research settings Chen et al. (2025), remains an underexplored but important direction.
@@
Given the limited bandwidth, our on-policy RL experiments are based on PPO algorithm and use 3B and 7B-scale models. Our observations and conclusions may be subject to change with more performant base models and different on-policy algorithms. This paper focuses on empirical experiments on public benchmarks. We believe this paper do not incur potential risks.
@@ Appendix A Dataset Statistics and Licenses
We report statistics of each dataset in Table˜4. Most datasets are released under permissive open licenses such as CC BY-SA (Natural Questions, SQuAD, HotpotQA, FEVER), CC BY (MS MARCO, DL19, DL20, SciFact), and Apache 2.0 (TriviaQA), while NFCorpus and FiQA restrict usage to academic or non-commercial purposes. SciFact uses ODC-By 1.0 for its corpus. ToolRet is a composite benchmark repurposed from multiple sources with mixed licensing; details are provided in Appendix B and Table 6 of Shi et al. (2025).
@@ Appendix B Results of SPQE with Different LLMs
Dense Contriever Retriever
@@
We report performance comparison of SPQE with different LLMs in Table˜5Table˜6 and Table˜7, respectively.
@@ Appendix C Reward Formulations
For retrieval reward, we report the specific instantiation for each task in Table˜8.

### v2

@@ 5.2 Experiment, Result and Analysis
For dense retriever, the benefits of our hybrid approach are most apparent. On the ad hoc retrieval results, the OPQE-7B model achieves the best average score of 58.1 in our setting, compared to the standard DR-7B (DeepRetrieval) model (57.5) and the prompt-based SPQE (56.6). This improvement is consistent across multiple datasets, with OPQE improving results on benchmarks like DL20 (70.5). Similarly, on the evidence-based retrieval datasets, the OPQE-3B model increases the average performance to 80.2, improving over other methods under the same retriever and training setup. This suggests that using RL to fine-tune pseudo-document generation can produce query representations that are effective for dense retrievers.
@@
For the sparse BM25 retriever, the results show that OPQE is highly competitive. While the zero-shot SPQE remains the top performer on evidence-based datasets (78.2), our OPQE-7B model closes the gap significantly and improves over the standard DR-7B model (76.5 vs. 75.6). On the ad hoc retrieval datasets, the OPQE-7B model matches the top performance of SPQE, with both achieving an average score of 51.6. In summary, by framing on-policy optimization as generating a pseudo-document rather than a rewritten query, our OPQE method combines the generative structure of prompting with the targeted fine-tuning of reinforcement learning, and yields strong performance across benchmarks.
Training Trajectories of OPQE.
@@
To better understand the performance gains of our hybrid OPQE method, we analyze its reward trajectories during training compared to the standard RL approach in Fig.˜1. These plots provide evidence for our hypotheses about the effectiveness of incorporating a pseudo-document generation structure.
@@
First, the trajectories validate our hypothesis that this structure effectively leverages the LLM’s internal knowledge, particularly for sparse retrievers. In Fig.˜1(a) and Fig.˜1(c) (FEVER and TriviaQA with BM25), the OPQE policy (blue line) starts with a significantly higher average reward than the standard RL policy (red line). This demonstrates that the initial, zero-shot pseudo-documents are already high-quality query expansions that provide rich lexical signals for the term-based BM25 retriever. OPQE essentially begins with a “warm start” by using the LLM’s knowledge, while the standard RL method must learn to generate effective query terms from a lower baseline.
@@ 6 Conclusion and Future Works
In this paper, we systematically evaluated and compared two LLM-based query augmentation paradigms: simple prompting-based query augmentation and RL-based query augmentation, across a variety of IR tasks. Our results suggest that prompt-based methods with advanced LLMs can perform on par with, or even surpass, RL-based methods in many scenarios, while RL-based approaches are particularly effective at enabling smaller LLMs to excel at this task. Furthermore, a hybrid approach, where pseudo-document generation prompts from prompt-based methods are used to provide stronger initialization for RL-based training, achieved the best overall performance. These findings suggest that RL-based query augmentation frameworks can be further optimized through improved task reformulation. Importantly, prompt-based methods like SPQE are directly usable in a zero-shot manner across datasets and retrievers, while on-policy RL approaches (including DeepRetrieval-style query rewriting and OPQE) typically require training a separate policy for each new corpus or task.
While our study highlights the strengths and trade-offs of prompting-based and RL-based query augmentation, several avenues remain open for future exploration. First, developing more lightweight yet effective prompting strategies for zero-shot query augmentation could reduce reliance on large LLMs and broaden applicability in resource-constrained settings. Second, refining RL task formulations, for example, by designing more nuanced reward functions, may help stabilize training and yield stronger performance. Finally, exploring query augmentation prompt design and RL training beyond the classic single-round retrieval paradigm, such as in multi-hop or deep research settings (Chen et al., 2025), remains an important direction.
@@
Our on-policy RL experiments focus on PPO and use lightweight policies (<7B), due to compute constraints. Our SPQE results are not backbone-matched to the DR-3B/7B policies: we use stronger instruction-tuned LLMs for SPQE to reflect the inference-time compute that prompting can leverage, and in preliminary experiments, SPQE with the same Qwen2.5-{3B,7B}-Instruct backbones was not as competitive. In our experiments, RL policies are trained separately per dataset, and we do not study cross-domain transfer following DeepRetrieval’s formulation; in practice, re-training a dedicated policy for each new corpus/task can be a bottleneck compared to zero-shot prompting baselines. Like other retrieval-optimized training setups, the RL results can be sensitive to reward design and prompt templates, and may require careful tuning when transferring to new retrievers, corpora, or evaluation metrics. Finally, while OPQE improves over direct RL query rewriting in our setting, OPQE’s formulation inherently changes the output structure and typical length distribution relative to direct rewriting, so attributing gains to a single factor (e.g., length alone) is not well-posed without introducing artificial constraints that change the method; we leave more targeted analyses to future work.
This work is based on public benchmarks, but tool retrieval benchmarks may contain overlapping or near-duplicate tools across sources; such overlap can amplify gains from query expansion and may not fully reflect deployment settings with frequently updated tool inventories. We do not introduce new user-facing systems, and we do not anticipate additional ethical risks.
@@ Appendix A Dataset Statistics and Licenses
We report statistics of each dataset in Table˜5. Most datasets are released under permissive open licenses such as CC BY-SA (Natural Questions, SQuAD, HotpotQA, FEVER), CC BY (MS MARCO, DL19, DL20, SciFact), and Apache 2.0 (TriviaQA), while NFCorpus and FiQA restrict usage to academic or non-commercial purposes. SciFact uses ODC-By 1.0 for its corpus. ToolRet is a composite benchmark repurposed from multiple sources with mixed licensing; details are provided in Appendix B and Table 6 of Shi et al. (2025).
@@ Appendix B Results of SPQE with Different LLMs
Dense Contriever-msmarco Retriever
@@
We report performance comparison of SPQE with different LLMs in Table˜6Table˜7 and Table˜8, respectively.
@@ Appendix C Reward Formulations
For retrieval reward, we report the specific instantiation for each task in Table˜9.